Suppose we want to determine the (binomial) probability (p) of getting 4 heads in 10 flips of a 2-sided coin. Using the Binomial Table in the appendix of the text, what values of n, x and p would we use to look up this probability, and what would be the probability?
Answer by greenestamps(13367)



Obviously we can't use the table in the appendix of your text to answer your question....!

Since the problem is about binomial probability, n is probably the number of trials (10), x is probably the number of favorable outcomes (heads; 4), and p is the probability of getting heads on each flip (presumably 1/2, since the coins should be fair coins if the statement of the problem doesn't say otherwise.)

To find the probability of getting 4 heads on 10 flips of a fair coin, without using your table, the denominator of the probability fraction is 2^10=1024, because there are two equally likely outcomes on each flip; the numerator is the number of ways 4 of the 10 flips can be heads, which is "10 choose 4", 10C4 = 210.

The probability is then 210/1024 = 105/512.